Is it possible for the average total cost curve to decline if the marginal cost curve is rising? Explain your answer.

Would be greatly appreciated if it answered in 5sentences by yout own not copy and paste
Answer #1

It is possible that average total cost will decline, if marginal cost curve is rising. It will happen as long as MC is less than ATC up to that particular level of output. Once MC for a unit is higher than the ATC before that unit of output, then ATC will also increase with increase in output. So, ATC will first decrease, then increase depending upon the rise in MC. It can be understood by the follow example.

Let,

Fixed cost = 10

Variable cost per unit till unit 4 = $4

So, ATC till unit 4 = (10 + 4*4)/4 = 6.5

Now, marginal cost for unit 5 = $5

Then,

ATC for 5 units = (10+4*4+5)/5 = $6.2

Now, marginal cost for unit 6 = $5.5

Then,

ATC for 6 units = (10+4*4+5+5.5)/6 = $6.08

Now it can be seen that MC is increasing from $4 to $5 then $5.5, but ATC is decreasing from $6.5 to $6.2 to $6.08.

But, When,

MC for unit 7 = $8

Then, ATC for 7 units = (10+4*4+5+5.5+ 8)/7 = $6.35

Now, ATC is also increasing along with MC.